OpenStudy (anonymous):

Solve the equation 2(3^x) = 3(4^x). Give an exact answer and a decimal approximation rounded to one decimal place.

OpenStudy (anonymous):

log(2) + log(3^x) = log(3) + log(4^x) = log(2) + x log(3) = log(3) + x log(4) = x log(3) - x log(4) = log(3) - log(2) = x(log(3)-log(4) / log(3) - log(4) = log(3) - log(2) / log(3) - log(4) The answer should be approximately 1.4 but I can't get the right answer

OpenStudy (anonymous):

I get -.7558684902 on my calculator

OpenStudy (anonymous):

Did I make an error in one of my steps?

OpenStudy (isaiah.feynman):

Not sure of the question.

OpenStudy (anonymous):

Well, I have to give an exact answer and a decimal approximation but I'm not sure I get two answers.

OpenStudy (anonymous):

My teacher has the answer as approximately - 1.4

OpenStudy (kropot72):

\[\frac{2}{3}\times3^{x}=4^{x}\] \[\ln \frac{2}{3}+x \ln 3=x \ln 4\] -0.4005 + 1.0986x = 1.3863x 0.2877x = -0.4055 x = -1.41

OpenStudy (anonymous):

so was I going about it the wrong way with all my log steps?

OpenStudy (kropot72):

Logs are needed, but just in one step.

OpenStudy (anonymous):

How did you get 2/3 instead of log(2)?

OpenStudy (kropot72):

Sorry, I missed out my first step. This was to divide both sides of the original equation by 3.

OpenStudy (anonymous):

2(3^x) / 3 = 3(4^x) / 3

OpenStudy (kropot72):

2(3^x) / 3 = 3(4^x) / 3 Which can be written as: \[\frac{2}{3}\times3^{x}=\frac{3}{3}\times4^{x}\]

OpenStudy (anonymous):

OK, I got a different answer when I took the log of 2/3

OpenStudy (anonymous):

ln(2/3) = -.4055 + 1.0986 = 1.3863 .6931x = 1.3863 1.3863 / .6931 = 2.0001

OpenStudy (anonymous):

I did something wrong because not ending up with 1.4

OpenStudy (anonymous):

How did you get -.4005 for ln(2/3)?

OpenStudy (anonymous):

It is coming up on my calculator as -.4054

OpenStudy (kropot72):

Sorry, it is bad copying on my part. I typed -0.4005 instead of -0.4055. However the rest of what I typed in my original post did not carry the error into calculation.

OpenStudy (anonymous):

So tomorrow for my final, when the teacher wants an exact answer and a decimal approximation, that just means not to round?

OpenStudy (anonymous):

-.4055 + 1.0986 = 1.3863 rounded to one decimal place is 1.4 .6931x = 1.3863, solve for x and that is exact answer?

OpenStudy (anonymous):

She only gave -1.4 as the answer, so I don't know what she means by an exact answer and a decimal approximation

OpenStudy (kropot72):

-0.4055 + 1.0986x = 1.3863x 0.2877x = -0.4055 x = -1.40945 (exact answer) x = -1.4 (rounded to one decimal place)

OpenStudy (anonymous):

Where does the - come from? I'm thinking -.4055 is positive

OpenStudy (anonymous):

Only because you are adding 1.0986 which is a bigger number

OpenStudy (kropot72):

-0.4055 + 1.0986x = 1.3863x Do you agree so far, that -0.4055 is valid being the natural log of 2/3?

OpenStudy (anonymous):

Yes, I agree with that

OpenStudy (kropot72):

-0.4055 + 1.0986x = 1.3863x Subtracting 1.0986x from both sides gives: -0.4055 = 1.3863x - 1.0986x = 0.2877x ...........(1) We can interchange the left and right sides of equation (1) to give: 0.2877x = -0.4055 Are you with me here?

OpenStudy (anonymous):

OK, I'm up to that point

OpenStudy (anonymous):

I did .2903 and not .2877

OpenStudy (anonymous):

I did get .2903 rather than .2877

OpenStudy (kropot72):

So do you agree that you made an error when subtracting: 1.3863x - 1.0986x

OpenStudy (anonymous):

I agree but now to find the error...

OpenStudy (anonymous):

Just found it, I entered it incorrectly

OpenStudy (kropot72):

np. We can all do that :)

OpenStudy (anonymous):

The final step is just to divide by .2877 on both sides

OpenStudy (anonymous):

x is approximately 1.409454293

OpenStudy (anonymous):

That should be -1.409454293

OpenStudy (kropot72):

Looks good. Perhaps taking it to five decimal places is enough for an 'exact' answer. Actually, strictly speaking, it is not possible to give an exact answer in such a calculation.

OpenStudy (anonymous):

I would just her an answer with all the decimal places on the calculator.

OpenStudy (anonymous):

give

OpenStudy (kropot72):

Hopefully that will get full marks.